### v1

@@ C.1 Significance of New GPT-4 Based Metric
GPT-4 prompt design For each of the three tasks, we designed three prompts for scoring. For Detailed Description Task (H2, C3, and M2), we designed the prompt in five areas: text length, text content variation, text detail variation, logical text description, and linguistic ability. For Counter-intuitiveness Reasoning Task (H3, C3, and M3), we designed the prompt in six areas: expressiveness of language, the logic of response, the common sense of response, understanding of counter-intuition, differences in text detail, and length of text. For Title Task (H4 and C4), we used the description, comprehension, and title of the manually annotated video as a reference to score the new title. Each Prompt can be seen in Figure 11, 12, and 13.
@@
Figure 11: GPT-4 prompt for Title Evaluation. We give GPT prompt as system, and then input a description, explanation, and two titles, the first one is our annotation, and the second one is the output of the model, and GPT-4 will evaluate the similarity between these two texts according to the prompt’s requirements.
@@
Figure 12: GPT-4 prompt for Description Evaluation. We give GPT-4 prompt as system, and then input two descriptions, the first one is our annotation, and the second one is the output of the model, and GPT-4 will evaluate the similarity between these two texts according to the prompt’s requirements.
@@
Figure 13: GPT-4 prompt for Explanation Evaluation. We give GPT prompt as system, and then input two explanations, the first one is our annotation, and the second one is the output of the model, and GPT-4 will evaluate the similarity between these two texts according to the prompt’s requirements.
@@
Instability of GPT-4 responses During the experimental process using GPT-4 as the metric, we found that in a few cases, GPT-4 would provide different evaluations for the same content. In such cases, we would perform multiple evaluations and take the average. Figure 14 shows an example of GPT-4 generating different responses to the same content.
Figure 14: GPT-4 generates different responses to the same content. User Input is the two sentences input to GPT-4, which is used to judge the similarity. The GPT-4 Output 1 and GPT-4 Output 2 below are the answers given by GPT-4 in two different conversations, and you can see that the two assessments are different.
C.2 On Failure of the Timestamp Localization Task
There is no model that has the ability to answer the Timestamp Localization Task. Figure 15 shows exactly what is answered for each model.
Figure 15: Model responses on Timestamp Localization Task. As can be seen, the model gives answers to the Timestamp Localization Task still focusing on describing the video content and still does not answer the specific time period when asked to type the number of frames and seconds.
C.3 Implementation Details
@@
All of our experiments took five to six days on a single NVIDIA A40 48GB GPU.
C.4 More Examples
Figure 16 shows the responses of different models on CreativeQA and MagicQA.
@@ C.4 More Examples
Figure 16: Model responses on CreativeQA and MagicQA. For the description of the Creative video example, only VideoChat gives the key point of the bicycle, but its description also has many errors and omissions, and the remaining two models do not identify the bicycle. In the explanation task, the responses of all three models fail to clearly explain the creativity of this imitation performance. For the Magic video example, all three models perform very poorly in description and explanation, basically only answering the phone and the straw, but lacking the description and explanation of the magic effect.

### v2

@@ 0.D.1 Significance of New GPT-4 Based Metric
GPT-4 prompt design For each of the three tasks, we designed three prompts for scoring. For Detailed Description Task (H2, C3, and M2), we designed the prompt in five areas: text length, text content variation, text detail variation, logical text description, and linguistic ability. For Counter-intuitiveness Reasoning Task (H3, C3, and M3), we designed the prompt in six areas: expressiveness of language, the logic of response, the common sense of response, understanding of counter-intuition, differences in text detail, and length of text. For Title Task (H4 and C4), we used the description, comprehension, and title of the manually annotated video as a reference to score the new title. Each Prompt can be seen in Fig. A11, A12, and A13.
@@
Figure A11: GPT-4 prompt for Title Evaluation. We give GPT prompt as system, and then input a description, explanation, and two titles, the first one is our annotation, and the second one is the output of the model, and GPT-4 will evaluate the similarity between these two texts according to the prompt’s requirements.
@@
Figure A12: GPT-4 prompt for Description Evaluation. We give GPT-4 prompt as system, and then input two descriptions, the first one is our annotation, and the second one is the output of the model, and GPT-4 will evaluate the similarity between these two texts according to the prompt’s requirements.
@@
Figure A13: GPT-4 prompt for Explanation Evaluation. We give GPT prompt as system, and then input two explanations, the first one is our annotation, and the second one is the output of the model, and GPT-4 will evaluate the similarity between these two texts according to the prompt’s requirements.
@@
WUPS The Wu-Palmer similarity (WPUS) measure calculates how similar two word senses are. It considers the depths of the two synsets in the WordNet taxonomies, along with the depth of the Least Common Subsumer (LCS). The WUPS metric has certain limitations that make it difficult to use in VQA tasks. First, certain words are very similar in vocabulary, but their meanings can be very different. The problem may arise with color. For example, if the answer to a certain question is white and the system predicts the answer to be black, that answer will still get a WUPS score of 0.92, which seems high. Another limitation is that WUPS cannot be used for answers to phrases or sentences because it always deals with rigid semantic concepts, which are most likely to be single words. More examples as shown in A14.
@@
Figure A14: The responses of VLMs to NExT-OE are scored using the WUPS and GPT-4 based metric. It can be observed that Otter’s responses, whether correct or close to the answer, receive scores near zero under the WUPS metric. In contrast, HGA’s completely incorrect responses (e.g., answering “cat” when the correct answer is “dog”) score highly under WUPS. On the other hand, the scoring by GPT-4 appears to be accurate.
Instability of GPT-4 responses During the experimental process using GPT-4 as the metric, we found that in a few cases, GPT-4 would provide different evaluations for the same content. In such cases, we would perform multiple evaluations and take the average. Fig. A15 shows an example of GPT-4 generating different responses to the same content.
Figure A15: GPT-4 generates different responses to the same content. User Input is the two sentences input to GPT-4, which is used to judge the similarity. The GPT-4 Output 1 and GPT-4 Output 2 below are the answers given by GPT-4 in two different conversations, and you can see that the two assessments are different.
0.D.2 On Failure of the Timestamp Localization Task
There is no model that has the ability to answer the Timestamp Localization Task. Fig. A16 shows exactly what is answered for each model.
Figure A16: Model responses on Timestamp Localization Task. As can be seen, the model gives answers to the Timestamp Localization Task still focusing on describing the video content and still does not answer the specific time period when asked to type the number of frames and seconds.
0.D.3 Implementation Details
@@
mPLUG-Owl mPLUG-Owl is a novel training paradigm that equips LLMs with multi-modal abilities through modularized learning of foundation LLM, a visual knowledge module, and a visual abstractor module. The training paradigm involves a two-stage method for aligning image and text, which learns visual knowledge with the assistance of LLM, while maintaining and even improving the generation abilities of LLM. In the second stage, languageonly and multi-modal supervised datasets are used to jointly fine-tune a low-rank adaption (LoRA) module on LLM and the abstractor module by freezing the visual knowledge module.
Video-LLaMA Video-LLaMA is built on top of BLIP-2 and MiniGPT-4. It is composed of two core components: (1) Vision-Language (VL) Branch (Visual encoder: ViT-G/14 + BLIP-2 Q-Former) and (2) Audio-Language (AL) Branch (Audio encoder: ImageBind-Huge). In VL Branch, a two-layer video Q-Former and a frame embedding layer (applied to the embeddings of each frame) are introduced to compute video representations. And In AL Branch, a two-layer audio Q-Former and an audio segment embedding layer (applied to the embedding of each audio segment) are introduced to compute audio representations.
0.D.4 Ablation Experiment about Audio
Audio is undoubtedly an important part of video, but existing models utilize audio in a very limited way, some of them don’t use audio as an input, and then get high scores on various lists. For example, Otter, mPLUG, video-ChatGPT and GIT in the paper, the outputs of these models are not affected by the audio, and only the visual information is understood; there are also some models that convert the audio into captions for text encoding and then input them into the model, such as VideoChat in the paper. However, The FunQA dataset is designed to emphasize the understanding of ‘fun’ through the visual aspect of videos. During the data collection phase, we focused on the visual elements, and in our experimental setup, we retained the audio of the videos when feeding them to the models. To investigate the role of audio, we conducted a set of ablation experiments by muting the audio during both training and testing phases. The results are shown in Table A5. This demonstrates that FunQA is an visual-centric dataset and audio could not ease the challenge of the benchmark.
Table A5: The ablation experiment about audio. The scores of the results after inputting the muted video, the traditional evaluation scores remain basically unchanged, and the scores of GPT-4 fluctuate very little (this is mainly due to the issues of instability of the evaluation method of GPT-4.
HumorQA CreativeQA MagicQA
H1
H2 H3 H4
C1
C2 C3 C4 C5
M1
M2 M3
VideoChat [36]
0.5 / 13.7 / 0.0 44.0 / 37.9 0.5 / 13.5 / 0.0 45.4 / 31.9 0.8 / 5.1 / 0.5 20.2 / 61.7
0.3 / 7.5 / 0.0 21.7 / 10.9 0.3 / 7.7 / 0.0 22.8 / 27.7 0.2 / 1.2 / 0.2 7.3 / 51.1 67.5
0.6 / 15.5 / 0.0 47.4 / 14.2 0.3 / 9.2 / 0.0 43.1 / 24.6
VideoChat (MUTED) [36]
0.5 / 13.5 / 0.0 40.0 / 34.5 0.5 / 12.5 / 0.0 43.7 / 31.9 0.8 / 5.1 / 0.5 22.4 / 57.8
0.3 / 7.4 / 0.0 24.0 / 10.0 0.4 / 7.7 / 0.0 20.2 / 30.5 0.2 / 1.2 / 0.2 6.8 / 45.3 67.5
0.6 / 15.0 / 0.0 48.0 / 12.2 0.3 / 9.2 / 0.0 44.0 / 23.4
Table A6: Human Performance on FunQA. 100 QA pairs (H2, H3, H4) from HumorQA in FunQA were randomly selected and answered by five different people. We use the following metrics: BLEU-4 / ROUGE-L / CIDEr / BLEURT / GPT-4 for evaluation.
Task H2 H3 H4
Human 1 6.0 / 28.8 / 6.5 / 46.6 / 82.4 2.2 / 20.8 / 1.1 / 45.0 / 85.2 0.8 / 9.2 / 1.1 / 20.5 / 76.6
Human 2 13.8 / 42.5 / 11.1 / 55.4 / 90.5 13.2 / 34.1 / 10.0 / 54.4 / 87.5 1.1 / 12.8 / 2.4 / 33.5 / 80.6
Human 3 0.7 / 16.5 / 0.5 / 36.2 / 77.5 0.7 / 14.7 / 0.2 / 44.1 / 69.0 0.2 / 3.1 / 0.3 / 15.4 / 76.8
Human 4 26.4 / 50.5 / 27.7 / 75.8 / 80.3 10.3 / 36.8 / 11.5 / 74.3 / 90.0 2.8 / 21.7 / 6.0 / 43.4 / 73.4
Human 5 3.5 / 16.7 / 3.7 / 48.0 / 86.6 2.4 / 16.4 / 1.8 / 47.9 / 90.3 0.8 / 10.5 / 0.8 / 23.9 / 88.2
Model (SOTA) 1.5 / 18.1 / 1.0 / 44.0 / 37.9 1.3 / 15.4 / 0.5 / 25.9 / 61.7 1.1 / 8.8 / 0.7 / 25.9 / 61.7
Human (Avg) 10.1 / 31.0 / 9.9 / 60.0 / 83.6 5.8 / 24.6 / 5.0 / 53.1 / 84.4 1.1 / 11.46 / 2.1 / 27.3 / 79.1
0.D.5 More Examples
We conducted experiments to compare machine versus human performance on FunQA. We randomly selected 100 HumorQA pairs from HumorQA and had five individuals who had not watched any FunQA videos provide answers to the questions (H2, H3, H4) for each video. Subsequently, we calculated the average score for each individual under the FunQA benchmark and the overall average score of human responses. The FunQA benchmark score is the selected SOTA score between models.The results are in Table A6.
0.D.6 More Examples
Fig. A17 shows the responses of different models on CreativeQA and MagicQA.
@@ 0.D.6 More Examples
Figure A17: Model responses on CreativeQA and MagicQA. For the description of the Creative video example, only VideoChat gives the key point of the bicycle, but its description also has many errors and omissions, and the remaining two models do not identify the bicycle. In the explanation task, the responses of all three models fail to clearly explain the creativity of this imitation performance. For the Magic video example, all three models perform very poorly in description and explanation, basically only answering the phone and the straw, but lacking the description and explanation of the magic effect.
Appendix 0.E More Discussions
0.E.1 (Potential) More Essential Factors
Accurate understanding of the videos Through our analysis of failure cases, we’ve observed that many models struggle with accurately describing videos. While they might be adept at detecting objects within the videos, they often falter in comprehending the contextual relationship between sequential events. Such misinterpretations indicate that there’s a need for further exploration in this domain. The videos we’ve used can indeed serve as an invaluable dataset for probing video descriptions in depth.
Logic Reasoning The primary nature of our videos encompasses content that is counterintuitive and contradicts common sense. For models to understand these, it’s imperative they grasp the concept of "common sense." They must deduce what would typically transpire under normal circumstances and then use that perspective to humorously interpret the video. This necessitates the model to possess strong reasoning capabilities, especially when it comes to common sense reasoning.
Extra Knowledge - Sense of Humor To decipher the humor in a video, it’s plausible that understanding the fundamental principles of humor is crucial. This type of knowledge, along with many other tidbits of common sense and additional information, might enhance the model’s performance. Determining how to integrate valuable knowledge and discerning what counts as "valuable" are topics that warrant further exploration.
0.E.2 Potential Solutions
Model Size Increasing the number of parameters is a natural method to enhance the model’s performance. However, this approach comes with its own set of engineering challenges, requiring improvements in model optimization and deployment. We’re also curious about the relationship between the number of parameters and the performance on FunQA. This is an intriguing research point in itself, and our dataset can serve as an excellent test bed to further this exploration.
Data Quality We believe the emphasis for this task should be on data collection. Current trends with large and dynamic models have shown that having vast amounts of low-quality data isn’t as effective as a smaller quantity of high-quality data. Thus, we hope the community can discover the type of data that genuinely assists in understanding counterintuitive videos. This is a crucial research direction.
Training Strategy Studying training strategies is also essential. For instance, determining which type of data to start learning from, and understanding the significance of curriculum learning, among others.
Model Collaboration Ultimately, we might not need to solely focus on a single model to solve this problem. Perhaps multiple models collaboratively working on examples in an elegant manner could be a method to enhance performance. However, this approach might necessitate paying more attention to the overall efficiency of model implementation
0.E.3 The Emphasis of Temporal Dynamics in FunQA
HumorQA Example: Humor_Example.mp4
H2: An individual slipped on the staircase filled with ice and tumbled down to the very bottom, followed by a second person who also fell after witnessing the first person’s fall.
H3: The first person falling down the stairs step by step was already very funny, and the second person repeating the same mistake and falling down made it even more hilarious.
Explanation: As answered in H2, the main element of humor in this video is two people slipping down an icy staircase in a sequential order, and if the model does not make sense of the temporal information, it will not be able to give the sequential logic of the two people slipping down one after the other. It is also clear from the answer to H3 that the sequence of slipping backwards and forwards is one of the sources of humor, as the person at the back does not learn from the lesson of the former and slips down in a similar way, and this repetition of the wrongdoing brings humor to the situation.
CreativeQA Example: Creative_Exapmle.mp4
C2: At the center of the stage is a blue rectangular box. Following the little girl’s watering of the box, four individuals lying inside gradually lift their hands and legs, each at different paces and heights, until they stop and reveal green painted leaf-shaped objects on their limbs.
C3: The main creative element of this video is when the four individuals in a box, with varying heights and movement speeds, gradually raise their hands and feet to mimic the growing process of the carrot seeds planted by the little girl. The green leaf-shaped objects, tied and opened in advance with their hands and feet, are used to simulate the true sprouting of carrot sprouts, resulting in a lively and imaginative scene.
Explanation: The performance in this video mimics growing bean sprouts, the process of growing beans into sprouts by constantly watering the soil. Trying to describe the content of this video idea and explain where the idea came from requires an understanding of the temporal information. Analyzing the visual information together, the model has to understand the sequence of watering, bean sprouts growing from the soil, and bean sprouts growing taller and greener in order to answer the question accurately. This difference in understanding can be demonstrated on the C3 task, where without analyzing the temporal information, the optimal answer would be bean sprouts growing from the soil, and cause and effect logical relationships such as watering to grow bean sprouts would be ignored.